i walk 20 meters down the hall to my class and i make it there in 8 seconds. what is my speed in meters per seconds?
Answer by jim_thompson5910(35256)

(20 meters)/(8 seconds) = 20/8 = 10/4 = 5/2 = 2.5

Your speed in m/s is 2.5 m/s